what comes next in the sequence.
1,3,7,13

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48558)

what comes next in the sequence.
1,3,7,13
next is 13+4 = 17
===========
Cheers,
Stan H.

Answer by checkley71(8403)

1,3,7,13,21
RULE:
1+(2)=3
3+(4)=7
7+(6)=13
13+(8)=21